A television costs $7 less than 4 times the cost of a microwave. The cost of both items is $438.
Answer by fractalier(6550)

Call the price of the microwave, x.
Then the price of the TV is 4x-7.
Together they add to 438, or
x + 4x-7 = 438
5x-7=438
5x=445
x=$89 for the microwave
4(89)-7= $349 for the TV